OpenStudy (anonymous):

Find a1 in the arithmetic sequence for which d = -3 and a17 = -38

OpenStudy (anonymous):

a17=a1+16d=-38 a1=-38-16d a1=-38-16(-3) a1=10

OpenStudy (anonymous):

a1=a=13 since a17=-83 then the formula is ar^16 it is 16 because they start at one but our formulas exponents start at 0 -83=13*r^16 sfr do you know how to solve for r?

OpenStudy (anonymous):

a=13 and a17=-83. => a+16d=-83 => d=-6 now put this in the formula for the sum S=35/2[2*13+34*-6]=-3115

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Arithmetic_progression

OpenStudy (anonymous):

there